OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

jimthompson5910 (jim_thompson5910):

y = mx+b -1 = 6(5) + b -1 = 30 + b -1 - 30 = b -31 = b b = -31 So the equation is y = 6x-31